# Fundamentals and Advances in Transverse Thermoelectrics

### v1

@@ Abstract
Transverse thermoelectric effects interconvert charge and heat currents in orthogonal directions due to the breaking of either time-reversal symmetry or structural symmetry, enabling simple and versatile thermal energy harvesting and solid-state cooling within single materials. In comparison to the complex module structures required for the conventional Seebeck and Peltier effects, the transverse thermoelectric effects provide the complete device structures, potentially resolving the fundamental issue of multi-module degradation of thermoelectric conversion performance. This review article provides an overview of all currently known transverse thermoelectric conversion phenomena and principles, as well as their characteristics, and reclassifies them in a unified manner. The performance of the transverse thermoelectric generator, refrigerator, and active cooler is formulated, showing that thermal boundary conditions play an essential role to discuss their behaviors. Examples of recent application research and material development in transverse thermoelectrics are also introduced, followed by a discussion of future prospects.
@@ I Introduction
The thermoelectric conversion performance of materials is often evaluated using the dimensionless figure of merit $zT$ with $T$ being the absolute temperature. As discussed in Sec. III in more detail, $zT$ for the Seebeck effect is proportional to the electrical conductivity and the square of the Seebeck coefficient and inversely proportional to the thermal conductivity. For a long time, $zT$ exceeding 1 was an indicator to realize practical applications of the thermoelectric conversion. As a result of active materials research in the 21st century, various materials with $zT$ far exceeding 1 have been synthesized and discovered [1]. However, despite the revolutionary progress in materials research, the range of applications of thermoelectric power generation and cooling technologies is still limited. Part of the reason for this situation is the complicated structure of the longitudinal thermoelectric modules based on the Seebeck and Peltier effects [Figs. 1(a) and 1(b)]. In this module structure, which has four junctions per pair of $p$-type and $n$-type conductors, the interfacial electrical and thermal resistances increase significantly; even if excellent thermoelectric characteristics are obtained at the material level, the energy conversion efficiency decreases when the multicomponent modules are constructed [1]. There are also problems with the thermal and mechanical durability of the modules and the manufacturing cost due to complexity.
@@
Figure 1: Schematics of (a) longitudinal thermoelectric generation, (b) longitudinal thermoelectric cooling, (c) transverse thermoelectric generation, and (d) transverse thermoelectric cooling. ${\bf J}$ and $\nabla T$ denotes the charge current and temperature gradient, respectively. The direction of ${\bf J}$ in (a) and (c) corresponds to the direction of the electric field driven by the thermoelectric effects, while that in (b) and (d) to the direction of the applied charge current.
@@
To quantify this situation, Fig. 2 compares the ideal reduced conversion efficiency $\bar{\eta}$ (red curve), i.e. the efficiency normalized by the Carnot efficiency, with the $\bar{\eta}$ values observed in actual longitudinal thermoelectric modules (blue plots) as a function of $ZT_{\rm ave}$, where $ZT_{\rm ave}$ is the averaged material $zT$ within the module measurement temperature range [2]. The $\bar{\eta}$ values refer to various thermoelectric modules, including Bi2Te3-based [3, 4, 5], PbTe-based [6], GeTe-based [7, 8], MgSi2-based [9], Mg3Sb2-based [10, 11], Skutterudite-based [12], and half-Heusler-alloy-based [13, 14] systems. In all cases, the observed efficiencies are much lower than the ideal values. Despite the development of thermoelectric modules with high $ZT_{\rm ave}$, the conversion efficiency has not increased drastically due to excess resistance, and the larger $ZT_{\rm ave}$, the greater the deviation from the ideal value. Following Refs. [8, 14, 15, 16], the resistance loss ratio is defined to be the ratio of the total excess resistance, including junctions and electrodes, to the total resistance of thermoelectric materials themselves. For simplicity, we assume that the resistance loss ratio is the same for both electrical and thermal resistances, and plot the $ZT_{\rm ave}$ dependence of $\bar{\eta}$ as the resistance loss ratio varies from 10% to 40% (Fig. 2). These curves agree well with the experimental data, indicating that high-performance longitudinal thermoelectric modules exhibit a resistance loss of several tens of percent.
@@
One way to solve these problems is to implement transverse thermoelectric conversion [17]. Using the transverse thermoelectric effects, it is possible to generate an electric voltage and current in a direction perpendicular to $\nabla T$, i.e. the transverse direction [Figs. 1(c) and 1(d)]. The material figure of merit $z_{xy}T$ for transverse thermoelectrics is defined in a similar way to that for longitudinal thermoelectrics; it is proportional to the electrical conductivity and the square of the transverse thermopower and inversely proportional to the thermal conductivity, where the transverse thermopower is defined as the generated electric field in the $y$ direction normalized by applied $\nabla T$ in the $x$ direction and the electrical (thermal) conductivity is for the $y$ ($x$) direction (see Secs. III and IV for details). The voltage and power induced by the transverse thermoelectric effects can be increased by increasing the length and area of the material in the direction perpendicular to $\nabla T$, respectively, and it is not necessary to form a large number of junction structures [Fig. 1(c)]. For this reason, the transverse thermoelectric conversion is suitable for reusing thermal energy distributed over a wide area. Furthermore, transverse thermoelectric conversion modules do not require junctions, and there are no problems of interfacial electrical and thermal resistances and thermal degradation on the hot side [note that two electrical contacts for extracting output can be attached to the cold side, as depicted in Fig. 1(c)]. Thus, the efficiency calculated from the material’s $z_{xy}T$ (red curve in Fig. 2) can be expected in experimental modules. The transverse thermoelectric conversion is also expected to improve the durability of the modules and reduce manufacturing costs. While the transverse thermoelectric conversion has many advantages, there are still various issues that remain to be addressed, such as the facts that the transverse thermopower has not yet reached a practical level and that the importance of thermal boundary conditions, which are essential for discussing the performance of the transverse thermoelectric devices, has not been fully recognized. Importantly, since various thermoelectric conversion principles are being studied independently in different fields, there is currently no unified paradigm for classifying these phenomena.
@@
In this article, we review fundamentals and advances of transverse thermoelectrics. First, in Sec. II, we systematically organize the classification of the transverse thermoelectric conversion phenomena and summarize their mechanisms and characteristics. The transverse thermoelectric conversion phenomena require the symmetry breaking for charge and heat carriers. Depending on how this symmetry is broken, the transverse thermoelectric conversion phenomena can be broadly classified into two categories: those that occur due to time-reversal symmetry breaking and those due to structural symmetry breaking. Here, the structural symmetry breaking in this article means that the sample has a certain structural asymmetry with respect to an axis taken in the direction of a charge or heat current, at either microsccale or macroscale. In other words, this terminology is defined as the breaking of sample’s spatial inversion symmetry about the left and right sides of that axis. In Secs. III and IV, we formulate the performance of a transverse thermoelectric generator, refrigerator, and active cooler for the cases with time-reversal symmetry breaking and structural symmetry breaking, respectively. Unlike the longitudinal thermoelectric conversion based on the Seebeck and Peltier effects, we show that thermal boundary conditions in the electric field direction play an essential role in the performance of the transverse thermoelectric conversion. For the transverse thermoelectric generator, the expressions of the figure of merit and efficiency vary depending on whether the thermal boundary conditions are isothermal or adiabatic, but the important point is that they are interchangeable with each other; the transverse thermoelectric generation is not inherently more or less efficient than the longitudinal thermoelectric generation, contrary to the discussion in Ref. [18]. Importantly, depending on the thermal boundary conditions, a correction term due to the transverse temperature gradient and Seebeck effect occurs in the transverse thermopower. The experiments shown in Sec. V demonstrate the importance of the contribution of this correction term. In Sec. VI, we summarize research activities for developing applications of the transverse thermoelectric conversion and introduce new concepts for further performance improvements. Section VII is devoted to the conclusions and prospects. We hope that this article serves as a cornerstone for basic and applied research on transverse thermoelectrics and provides an opportunity for interdisciplinary fusion of research on various transverse thermoelectric conversion phenomena that have been studied in different communities.
@@ II Classification of transverse thermoelectric conversion phenomena
The purpose of this section is to establish a unified classification for various transverse thermoelectric conversion phenomena. The transverse thermoelectric conversion phenomena caused by time-reversal symmetry breaking include the magneto-thermoelectric and thermo-spin effects, which have been studied mainly in the field of spin caloritronics [19, 20, 21]. Here, the magneto-thermoelectric effects refer to the thermoelectric effects depending on magnetic fields or magnetization, while the thermo-spin effects refer to the conversion between a heat current and a spin current, i.e. a flow of spin angular momentum. Some of these phenomena were discovered in the 21st century, and they are still in their infancy. Although the transverse thermoelectric conversion phenomena caused by structural symmetry breaking have been studied in the field of thermoelectrics for a long time [22, 23], they are often called by completely different terminologies, which may confuse non-specialists. For example, the transverse thermoelectric conversion in a material having charge carriers with opposite signs along different crystal orientations, i.e. axis-dependent conduction polarity (ADCP), is called goniopolarity [24, 25, 26] when it is caused by a single band, but called ($p\times n$)-type transverse thermoelectrics [27, 28] when it is caused by multiple bands, despite the fact that the details of the band structure cannot be determined only by thermoelectric measurements. We argue that the leading classification of the transverse thermoelectric effects should be based on the symmetry and phenomenology, which can be determined from thermoelectric measurements alone. Therefore, we propose a classification of the transverse thermoelectric effects comprising three hierarchies: symmetry (hierarchy 1), phenomenology (hierarchy 2), and mechanism (hierarchy 3). Figure 3 shows the currently known transverse thermoelectric generation phenomena organized based on this classification, which is a refinement of the classification in Ref. [17]. In the following subsections, we outline the mechanisms and characteristics of each phenomenon. Although we will mainly focus on the thermoelectric generation, the phenomena in Fig. 3 have their Onsager reciprocals, enabling transverse thermoelectric cooling (Table 1). Recently, a higher-order transverse thermoelectric effect called the transverse Thomson effect [29] has been observed experimentally, but it is not discussed in this article.
@@
Figure 3: Classification of the transverse thermoelectric effects and schematics of (a) the ordinary Nernst effect (ONE), (b) anomalous Nernst effect (ANE), (c) Seebeck-driven ordinary Hall effect, (d) Seebeck-driven anomalous Hall effect (SAHE), (e) spin-Seebeck-driven inverse spin Hall effect (ISHE), (f) spin-dependent-Seebeck-driven ISHE, (g) off-diagonal Seebeck effect due to macroscale anisotropic structure, and (h) off-diagonal Seebeck effect due to microscale anisotropic structure. ${\bf H}$, ${\bf M}$, ${\bf J}_{\rm s}$, $e^{-}$, and $h^{+}$ denote the magnetic field, magnetization, spatial direction of a spin current, conduction electrons, and holes, respectively. In (g), the gray (orange) arrows show the direction of charge (heat) currents in the constituent conductors of the artificially tilted multilayer (ATML).
@@
Seebeck-driven ordinary Hall effect Ordinary-Hall-driven Peltier effect
Seebeck-driven anomalous Hall effect Anomalous-Hall-driven Peltier effect
Spin-Seebeck-driven inverse spin Hall effect Spin-Hall-driven spin Peltier effect
Spin-dependent-Seebeck-driven inverse spin Hall effect Spin-Hall-driven spin-dependent Peltier effect
@@ II.1 Transverse thermoelectric effects with time-reversal symmetry breaking
ONE is caused by the Lorentz force acting on conduction electrons and/or holes in a conductor under a magnetic field [Fig. 3(a)]. The ONE-induced transverse thermopower in typical metals is very small, but it is known that several Dirac and Weyl semimetals exhibit high transverse thermoelectric conversion performance due to ONE [31, 32]. For example, the figure of merit of BiSb alloys reaches $z_{xy}T>0.3$ at 100–200 K when an external magnetic field of $\sim$1 T is applied [31]. However, the drawback of ONE is that an external magnetic field must be applied to generate the transverse thermopower. To overcome this drawback, it has been demonstrated that the thermoelectric power generation based on ONE can be achieved without applying an external magnetic field by embedding permanent magnets in a module, where the remanent magnetization of the permanent magnets applies stray magnetic fields to the adjacent conductors [33]. However, there are still issues to be addressed, such as the fact that the fill factor of ONE materials that are responsible for the power generation decreases by the amount of the embedded permanent magnets.
@@
ANE is a phenomenon in which a thermoelectric voltage is generated in the direction of the cross product of $\nabla T$ and the spontaneous magnetization ${\bf M}$ in a magnetic material. The transverse thermopower $S_{\rm T}$ in a magnetic material under a magnetic field is phenomenologically described as the summation of the contributions proportional to $H$, i.e. ONE, and magnetization $M$, i.e. ANE [34]:
@@
To realize practical applications of ANE, it is necessary to find and develop magnetic materials with larger $S_{\rm ANE}$. Guidelines for designing $S_{\rm ANE}$ can be obtained by separating it into two components:
@@
II.1.3 Seebeck-driven ordinary and anomalous Hall effects.
As part of efforts to further improve magneto-thermoelectric conversion performance, a transverse thermoelectric conversion mechanism appearing in hybrid materials composed of magnetic and thermoelectric materials was proposed and demonstrated in 2021 [63]. This mechanism, originally named the Seebeck-driven transverse thermoelectric generation, operates when a closed circuit is constructed by connecting the ends of the magnetic and thermoelectric materials and is driven by the concerted action of the Seebeck effect in the thermoelectric material and AHE in the magnetic material. In other words, this mechanism corresponds to artificially modulating the second term on the right-hand side of Eq. (2) by constructing hybrid materials. However, the name “Seebeck-driven transverse thermoelectric generation” could be confused with the transverse thermoelectric conversion due to the off-diagonal Seebeck effect described below. In this article, we thus call this mechanism the Seebeck-driven AHE (SAHE).
SAHE has a high degree of freedom in material design and the transverse thermopower in the hybrid structure depicted in Fig. 3(d) is expressed as
@@
where $\rho_{\rm M(TE)}$ and $S_{\rm M(TE)}$ are the longitudinal resistivity and Seebeck coefficient of the magnetic (thermoelectric) material, respectively, and $r$ is the size ratio determined by the dimensions of the magnetic and thermoelectric materials [63]. When the in-plane areas of the magnetic and thermoelectric materials are the same, $r$ is determined by the thickness ratio of the two: $r=t_{\rm TE}/t_{\rm M}$ with $t_{\rm M(TE)}$ being the thickness of the magnetic (thermoelectric) material. The second term on the right-hand side of Eq. (3) represents the contribution of SAHE. By optimizing the combination and size ratio of the magnetic and thermoelectric materials, it is possible to obtain a much larger transverse thermopower than $S_{\rm ANE}$ alone. When a thin film of a magnetic material is combined with a slab of a thermoelectric material, $r$ becomes very large, and $S_{\rm T}$ approaching $100~{}\mu{\rm V/K}$ has been observed [63]. When a slab of a magnetic material is combined with a slab of a thermoelectric material, $S_{\rm T}$ is limited to around 10-$20~{}\mu{\rm V/K}$ due to smaller $r$ but such all-bulk hybrid structure is more suitable for increasing output power [64]. Recently, direct-contact SAHE has also been demonstrated, in which a magnetic material is directly deposited on the surface of a thermoelectric material, rather than forming a closed circuit [65]. Since there is a vast amount of research data accumulated over many years on AHE and the Seebeck effect, there is still plenty of room for improving the performance of SAHE. However, its versatility is lower than that of ANE; whereas ANE works in both in-plane and perpendicularly magnetized configurations, SAHE works only in the perpendicularly magnetized configuration. Furthermore, even if $S_{\rm T}$ is enhanced, $z_{xy}T$ and power generation efficiency of SAHE are not always higher than those of ANE [66].
It is worth mentioning that, if a magnetic material showing AHE is replaced with a nonmagnetic material showing the ordinary Hall effect, the Seebeck-driven ordinary Hall effect should occur [Fig. 3(c)]. In principle, the transverse thermoelectric conversion due to the Seebeck-driven ordinary Hall effect exists in any hybrid material comprising two or more materials under a magnetic field. However, at present, there are no reports where its contribution is explicitly observed.

### v2

@@ Abstract
Transverse thermoelectric effects interconvert charge and heat currents in orthogonal directions due to the breaking of either time-reversal symmetry or structural symmetry, enabling simple and versatile thermal energy harvesting and solid-state cooling/heating within single materials. In comparison to the complex module structures required for the conventional Seebeck and Peltier effects, the transverse thermoelectric effects provide the complete device structures, potentially resolving the fundamental issue of multi-module degradation of thermoelectric conversion performance. This review article provides an overview of all currently known transverse thermoelectric conversion phenomena and principles, as well as their characteristics, and reclassifies them in a unified manner. The performance of the transverse thermoelectric generator, refrigerator, and active cooler is formulated, showing that thermal boundary conditions play an essential role to discuss their behaviors. Examples of recent application research and material development in transverse thermoelectrics are also introduced, followed by a discussion of future prospects.
@@ I Introduction
The thermoelectric conversion performance of materials is often evaluated using the dimensionless figure of merit $zT$ with $T$ being the absolute temperature. As discussed in Sec. III in more detail, $zT$ for the Seebeck effect is proportional to the electrical conductivity and the square of the Seebeck coefficient and inversely proportional to the thermal conductivity. For a long time, $zT$ exceeding 1 was an indicator to realize practical applications of the thermoelectric conversion. As a result of active materials research in the 21st century, various materials with $zT$ far exceeding 1 have been synthesized and discovered [1]. However, despite the revolutionary progress in materials research, the range of applications of thermoelectric power generation and cooling/heating technologies is still limited. Part of the reason for this situation is the complicated structure of the longitudinal thermoelectric modules based on the Seebeck and Peltier effects [Figs. 1(a) and 1(b)]. In this module structure, which has four junctions per pair of $p$-type and $n$-type conductors, the interfacial electrical and thermal resistances increase significantly; even if excellent thermoelectric characteristics are obtained at the material level, the energy conversion efficiency decreases when the multicomponent modules are constructed [1]. There are also problems with the thermal and mechanical durability of the modules and the manufacturing cost due to complexity.
@@
Figure 1: Schematics of (a) longitudinal thermoelectric generation, (b) longitudinal thermoelectric cooling/heating, (c) transverse thermoelectric generation, and (d) transverse thermoelectric cooling/heating. ${\bf J}$ ($\parallel y$) and $\nabla T$ ($\parallel x$) denotes the charge current and temperature gradient, respectively. The direction of ${\bf J}$ in (a) and (c) corresponds to the direction of the electric field driven by the thermoelectric effects. In (b) and (d), the direction of ${\bf J}$ is shown for the cooling operation.
@@
To quantify this situation, Fig. 2 compares the ideal reduced conversion efficiency $\bar{\eta}$ (red curve), i.e. the efficiency normalized by the Carnot efficiency, with the $\bar{\eta}$ values observed in actual longitudinal thermoelectric modules (blue plots) as a function of the device figure of merit $ZT_{\rm ave}$, where $ZT_{\rm ave}$ is defined as the averaged material $zT$ within the module measurement temperature range [2] (note that lowercase $z$ and uppercase $Z$ represent the material and device figures of merit, respectively). The $\bar{\eta}$ values refer to various thermoelectric modules, including Bi2Te3-based [3, 4, 5], PbTe-based [6], GeTe-based [7, 8], MgSi2-based [9], Mg3Sb2-based [10, 11], Skutterudite-based [12], and half-Heusler-alloy-based [13, 14] systems. In all cases, the observed efficiencies are much lower than the ideal values. Despite the development of thermoelectric modules with high $ZT_{\rm ave}$, the conversion efficiency has not increased drastically due to excess resistance, and the larger $ZT_{\rm ave}$, the greater the deviation from the ideal value. Following Refs. [8, 14, 15, 16], the resistance loss ratio is defined to be the ratio of the total excess resistance, including junctions and electrodes, to the total resistance of thermoelectric materials themselves. For simplicity, we assume that the resistance loss ratio is the same for both electrical and thermal resistances, and plot the $ZT_{\rm ave}$ dependence of $\bar{\eta}$ as the resistance loss ratio varies from 10% to 40% (Fig. 2). These curves agree well with the experimental data, indicating that high-performance longitudinal thermoelectric modules exhibit a resistance loss of several tens of percent.
@@
One way to solve these problems is to implement transverse thermoelectric conversion [17]. Using the transverse thermoelectric effects, it is possible to generate an electric voltage and current in a direction perpendicular to $\nabla T$, i.e. the transverse direction [Figs. 1(c) and 1(d)]. The material figure of merit $z_{xy}T=S_{yx}^{2}T/\rho_{yy}\kappa_{xx}$ for transverse thermoelectrics is defined in a similar way to that for longitudinal thermoelectrics, where the transverse thermopower $S_{yx}$ is defined as the generated electric field in the $y$ direction normalized by applied $\nabla T$ in the $x$ direction (see Secs. III and IV for details). Here, $\kappa_{xx}$ is the thermal conductivity for the $x$ direction and $\rho_{yy}$ is the electrical resistivity for the $y$ direction. The voltage and power induced by the transverse thermoelectric effects can be increased by increasing the length and area of the material in the direction perpendicular to $\nabla T$, respectively, and it is not necessary to form a large number of junction structures [Fig. 1(c)]. For this reason, the transverse thermoelectric conversion is suitable for reusing thermal energy distributed over a wide area. Furthermore, transverse thermoelectric conversion modules do not require junctions, and there are no problems of interfacial electrical and thermal resistances and thermal degradation on the hot side [note that two electrical contacts for extracting output can be attached to the cold side, as depicted in Fig. 1(c)]. Thus, the efficiency calculated from the material’s $z_{xy}T$ (red curve in Fig. 2) can be expected in experimental modules. The transverse thermoelectric conversion is also expected to improve the durability of the modules and reduce manufacturing costs. While the transverse thermoelectric conversion has many advantages, there are still various issues that remain to be addressed, such as the facts that the transverse thermopower has not yet reached a practical level and that the importance of thermal boundary conditions, which are essential for discussing the performance of the transverse thermoelectric devices, has not been fully recognized. Importantly, since various thermoelectric conversion principles are being studied independently in different fields, there is currently no unified paradigm for classifying these phenomena.
@@
In this article, we review fundamentals and advances of transverse thermoelectrics. This article is organized as follows. First, in Sec. II, we systematically organize the classification of the transverse thermoelectric conversion phenomena and summarize their mechanisms and characteristics. The transverse thermoelectric conversion phenomena require the symmetry breaking for charge and heat carriers. Depending on how this symmetry is broken, the transverse thermoelectric conversion phenomena can be broadly classified into two categories: those that occur due to time-reversal symmetry breaking and those due to structural symmetry breaking. Here, the structural symmetry breaking in this article means that the sample has a certain structural asymmetry with respect to an axis taken in the direction of a charge or heat current, at either microscale or macroscale. In other words, this terminology is defined as the breaking of sample’s spatial inversion symmetry about the left and right sides of that axis. In Secs. III and IV, we formulate the performance of a transverse thermoelectric generator, refrigerator, and active cooler for the cases with time-reversal symmetry breaking and structural symmetry breaking, respectively. Although a thermoelectric device also acts as a heat pump, this heating mode is not dealt with in this article due to the limited space. Unlike the longitudinal thermoelectric conversion based on the Seebeck and Peltier effects, we show that thermal boundary conditions in the electric field direction play an essential role in the performance of the transverse thermoelectric conversion. For the transverse thermoelectric generator, the expressions of the figure of merit and efficiency vary depending on whether the thermal boundary conditions are isothermal or adiabatic, but the important point is that they are interchangeable with each other; the transverse thermoelectric generation is not inherently more or less efficient than the longitudinal thermoelectric generation, contrary to the discussion in Ref. [18]. Importantly, depending on the thermal boundary conditions, a correction term due to the transverse temperature gradient and Seebeck effect occurs in the transverse thermopower. The experiments shown in Sec. V demonstrate the importance of the thermal boundary condition and the contribution of the correction term. In Sec. VI, we summarize research activities for developing applications of the transverse thermoelectric conversion and introduce new concepts for further performance improvements. Section VII is devoted to the conclusions and prospects. We hope that this article serves as a cornerstone for basic and applied research on transverse thermoelectrics and provides an opportunity for interdisciplinary fusion of research on various transverse thermoelectric conversion phenomena that have been studied in different communities.
@@ II Classification of transverse thermoelectric conversion phenomena
The purpose of this section is to establish a unified classification for various transverse thermoelectric conversion phenomena. The transverse thermoelectric conversion phenomena caused by time-reversal symmetry breaking include the magneto-thermoelectric and thermo-spin effects, which have been studied mainly in the field of spin caloritronics [19, 20, 21]. Here, the magneto-thermoelectric effects refer to the thermoelectric effects depending on magnetic fields or magnetization, while the thermo-spin effects refer to the conversion between a heat current and a spin current, i.e. a flow of spin angular momentum. Some of these phenomena were discovered in the 21st century, and they are still in their infancy. Although the transverse thermoelectric conversion phenomena caused by structural symmetry breaking have been studied in the field of thermoelectrics for a long time [22, 23], they are often called by completely different terminologies, which may confuse non-specialists. For example, the transverse thermoelectric conversion in a material having charge carriers with opposite signs along different crystal orientations, i.e. axis-dependent conduction polarity (ADCP), is called goniopolarity [24, 25, 26] when it is caused by a single band, but called ($p\times n$)-type transverse thermoelectrics [27, 28] when it is caused by multiple bands, despite the fact that the details of the band structure cannot be determined only by thermoelectric measurements. We argue that the leading classification of the transverse thermoelectric effects should be based on the symmetry and phenomenology, which can be determined from thermoelectric measurements alone. Therefore, we propose a classification of the transverse thermoelectric effects comprising three hierarchies: symmetry (hierarchy 1), phenomenology (hierarchy 2), and mechanism (hierarchy 3). Figure 3 shows the currently known transverse thermoelectric generation phenomena organized based on this classification, which is a refinement of the classification in Ref. [17]. In the following subsections, we outline the mechanisms and characteristics of each phenomenon. Although we will mainly focus on the thermoelectric generation, the phenomena in Fig. 3 have their Onsager reciprocals, enabling transverse thermoelectric cooling/heating (Table 1). Recently, a higher-order transverse thermoelectric effect called the transverse Thomson effect [29] has been observed experimentally, but it is not discussed in this article.
@@
Figure 3: Classification of the transverse thermoelectric effects and schematics of (a) the ordinary Nernst effect (ONE), (b) anomalous Nernst effect (ANE), (c) Seebeck-effect-driven ordinary Hall effect, (d) Seebeck-effect-driven anomalous Hall effect (SAHE), (e) spin-Seebeck-effect-driven (SSE-driven) inverse spin Hall effect (ISHE), (f) spin-dependent-Seebeck-effect-driven (SdSE-driven) ISHE, (g) off-diagonal Seebeck effect due to macroscale anisotropic structure, and (h) off-diagonal Seebeck effect due to microscale anisotropic structure. ${\bf H}$, ${\bf M}$, ${\bf J}_{\rm s}$, $e^{-}$, and $h^{+}$ denote the magnetic field, magnetization, spatial direction of a spin current, conduction electrons, and holes, respectively. In (g), the gray (orange) arrows show the direction of charge (heat) currents in the constituent conductors of the artificially tilted multilayer (ATML). The thermopower tensor is antisymmetric $S_{xy}=-S_{yx}$ when a magnetic field or magnetism breaks time-reversal symmetry, or symmetric $S_{xy}=S_{yx}$ when a crystal lattice or layered macrostructure breaks structural symmetry.
@@
Seebeck-effect-driven ordinary Hall effect Ordinary-Hall-effect-driven Peltier effect
Seebeck-effect-driven anomalous Hall effect Anomalous-Hall-effect-driven Peltier effect
Spin-Seebeck-effect-driven inverse spin Hall effect Spin-Hall-effect-driven spin Peltier effect
Spin-dependent-Seebeck-effect-driven inverse spin Hall effect Spin-Hall-effect-driven spin-dependent Peltier effect
@@ II.1 Transverse thermoelectric effects with time-reversal symmetry breaking
ONE is caused by the Lorentz force acting on conduction electrons and/or holes in a conductor under a magnetic field [Fig. 3(a)]. The ONE-induced transverse thermopower in typical metals is very small, but it is known that several Dirac/Weyl semimetals and low-dimensional materials exhibit high transverse thermoelectric conversion performance due to ONE [31, 32, 33, 34, 35]. For example, the figure of merit of BiSb alloys reaches $z_{xy}T>0.3$ at 100–200 K when an external magnetic field of $\sim$1 T is applied [31]. However, the drawback of ONE is that an external magnetic field must be applied to generate the transverse thermopower. To overcome this drawback, it has been demonstrated that the thermoelectric power generation based on ONE can be achieved without applying an external magnetic field by embedding permanent magnets in a module, where the remanent magnetization of the permanent magnets applies stray magnetic fields to the adjacent conductors [36]. However, there are still issues to be addressed, such as the fact that the fill factor of ONE materials that are responsible for the power generation decreases by the amount of the embedded permanent magnets.
@@
ANE is a phenomenon in which a thermoelectric voltage is generated in the direction of the cross product of $\nabla T$ and the spontaneous magnetization ${\bf M}$ in a magnetic material [Fig. 3(b)]. The transverse thermopower $S_{\rm T}$ in a magnetic material under a magnetic field is phenomenologically described as the summation of the contributions proportional to $H$, i.e. ONE, and magnetization $M$, i.e. ANE [37]:
@@
To realize practical applications of ANE, it is necessary to find and develop magnetic materials with larger $S_{\rm ANE}$. Guidelines for designing $S_{\rm ANE}$ can be obtained by separating it into two components [44, 65]:
@@
Here, we should mention the non-relativistic counterpart of ANE: the topological Nernst effect [Fig. 3(b)]. In the breaking of the spin-rotation invariance, the Nernst effect occurs in the absence of the spin-orbit interaction due to non-coplanar spin structures, e.g. in skyrmions and non-coplanar antiferromagnets [67, 68, 69, 70]. Although only a limited number of materials exhibit the topological Nernst effect, this phenomenon also functions as the transverse thermoelectric conversion.
II.1.3 Seebeck-effect-driven ordinary and anomalous Hall effects.
As part of efforts to further improve magneto-thermoelectric conversion performance, a transverse thermoelectric conversion mechanism appearing in hybrid materials composed of magnetic and thermoelectric materials was proposed and demonstrated in 2021 [71]. This mechanism, originally named the Seebeck-driven transverse thermoelectric generation, operates when a closed circuit is constructed by connecting the ends of the magnetic and thermoelectric materials and is driven by the concerted action of the Seebeck effect in the thermoelectric material and AHE in the magnetic material. In other words, this mechanism corresponds to artificially modulating the second term on the right-hand side of Eq. (2) by constructing hybrid materials. However, the name “Seebeck-driven transverse thermoelectric generation” could be confused with the transverse thermoelectric conversion due to the off-diagonal Seebeck effect described below. In this article, we thus call this mechanism the Seebeck-effect-driven AHE (SAHE).
SAHE has a high degree of freedom in material design. The transverse thermopower in the hybrid structure depicted in Fig. 3(d) is expressed as
@@
[…]
It is worth mentioning that, if a magnetic material showing AHE is replaced with a nonmagnetic material showing the ordinary Hall effect, the Seebeck-effect-driven ordinary Hall effect should occur [Fig. 3(c)]. In principle, the transverse thermoelectric conversion due to the Seebeck-effect-driven ordinary Hall effect exists in any hybrid material comprising two or more materials under a magnetic field. However, at present, there are no reports where its contribution is explicitly observed.